# Just Trial Once: Ongoing Causal Validation of Machine Learning Models

### v1

@@ Appendix B Proofs
    $f_{Y}^{L}(x,a,m,\epsilon_{Y})$ =$=\begin{cases}Y_{\text{min}},&\ \text{if }\mathcal{S}_{a}(x)=\varnothing,\\ Y_{\text{min}},&\ \text{if }\mathcal{S}_{a}(x)\neq\varnothing,m<\min(\mathbf{M}_{a}(x))\\ F_{Y\mid x,a,h_{\text{prev}}(m)}^{-1}(\epsilon_{Y}),&\ \text{if }\mathcal{S}_{a}(x)\neq\varnothing,m>\min(\mathbf{M}_{a}(x)),m\not\in\mathbf{M}_{a}(x),\\ F_{Y\mid x,a,m}^{-1}(\epsilon_{Y}),&\ \text{if }\mathcal{S}_{a}(x)\neq\varnothing,m\in\mathbf{M}_{a}(x),\end{cases}$ (57)
    $f_{Y}^{U}(x,a,m,\epsilon_{Y})$ =$=\begin{cases}Y_{\text{max}},&\ \text{if }\mathcal{S}_{a}(x)=\varnothing,\\ Y_{\text{max}},&\ \text{if }\mathcal{S}_{a}(x)\neq\varnothing,m>\max(\mathbf{M}_{a}(x))\\ F_{Y\mid x,a,h_{\text{next}}(m)}^{-1}(\epsilon_{Y}),&\ \text{if }\mathcal{S}_{a}(x)\neq\varnothing,m<\max(\mathbf{M}_{a}(x)),m\not\in\mathbf{M}_{a}(x),\\ F_{Y\mid x,a,m}^{-1}(\epsilon_{Y}),&\ \text{if }\mathcal{S}_{a}(x)\neq\varnothing,m\in\mathbf{M}_{a}(x),\end{cases}$ (58)
    where $F^{-1}_{Y\mid x,a,m}$ is the inverse conditional CDF of $Y$ given $X,A,M$, derived from $P$, and where $\mathcal{S}_{a}(x)\coloneqq\{i\in\mathcal{D}\mid\pi_{i}(x)=a\}$ and $\mathbf{M}_{a}(x)\coloneqq\{f_{M}(\pi_{i}):i\in\mathcal{S}_{a}(x)\}$, as defined previously above.
@@
We have now defined the SCMs $\mathcal{M}_{L},\mathcal{M}_{U}$, and shown that these SCMs are both consistent with our assumptions. We will now briefly verify that both SCMs give rise to the same observed distribution $P(X,A,M,Y,D)$, and then demonstrate that these SCMs achieve the upper and lower bounds that are given in Assumption 3.3 when using $\tilde{\mathcal{S}}$ in place of $\mathcal{S}$.
First, we have it by construction that both SCMs yield the observed distribution $P(X,A,M,D)$, so it remains to demonstrate that they agree with the observed distribution $P(Y\mid X,A,M,D)$, which we can write equivalently as $P(Y\mid X,A,M)$, since $D\perp\!\!\!\perp Y\mid X,A,M$ under our assumed data-generating process. Note that $P(Y\mid X,A,M)$ is only well-defined for $x,a,m$ with positive density (if $X$ is continuous) or probability mass (if $X$ is discrete). Assuming that $p(x)>0$ for all $x\in\mathcal{X}$, we have constructed $f_{Y}^{L},f_{Y}^{U}$ to agree with $P(Y\mid x,a,m)$ for all $a,m$ where there exists a trialed policy $\pi$ that outputs $a=\pi(x)$ with performance $m=f_{M}(\pi)$. We note that for any other value of $a^{\prime},m^{\prime}$, we have it that $p(a^{\prime},m^{\prime}\mid x)=0$, and hence the entire set $(x,a^{\prime},m^{\prime})$ has zero density, and it is precisely on these never-observed subsets of inputs where $\mathcal{M}_{L},\mathcal{M}_{U}$ disagree.
@@
$\mathbb{E}_{\mathcal{M}_{L}}[Y(A=\pi_{e},M=f_{M}(\pi_{e}))]$ $=\mathbb{E}\big{[}\mathbf{1}\{X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\}f_{Y}^{L}(X,\pi_{e}(X),f_{M}(\pi_{e}),\epsilon_{Y})$ (59)
$\qquad+\mathbf{1}\{X\in\mathcal{X}_{<},\pi_{e}(X)\neq a_{0}\}f_{Y}^{L}(X,\pi_{e}(X),f_{M}(\pi_{e}),\epsilon_{Y})$ (60)
$\qquad+\mathbf{1}\{X\in\mathcal{X}_{\not\leq},\pi_{e}(X)\neq a_{0}\}f_{Y}^{L}(X,\pi_{e}(X),f_{M}(\pi_{e}),\epsilon_{Y})$ (61)
$\qquad+\mathbf{1}\{X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}f_{Y}^{L}(X,\pi_{e}(X),f_{M}(\pi_{e}),\epsilon_{Y})$ (62)
$\qquad+\mathbf{1}\{X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}f_{Y}^{L}(X,\pi_{e}(X),f_{M}(\pi_{e}),\epsilon_{Y})]$ (63)
@@
  • For Eq. 59, we can observe that for all $x\in\mathcal{X}_{=},a=\pi_{e}(x)$, the set $\mathcal{S}_{a}(x)$ is non-empty by definition of $\mathcal{X}_{=}$, and moreover that $m_{e}\in\mathbf{M}_{a}(x)$, where we define $m_{e}\coloneqq f_{M}(\pi_{e})$. In this case, we have it that $f_{Y}^{L}(x,a,m_{e},\epsilon_{Y})=F^{-1}_{Y\mid x,a,m_{e}}(\epsilon_{Y})\sim P(Y\mid x,a,m_{e})$, and so the term Eq. 59 is equal to $\mathbf{1}\{X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e}(X),M=f_{M}(\pi_{e})]$, which is equal to
    $\mathbf{1}\{X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{=}(X)]$ (64)
    from Eq. 18 of Section B.1.
  • For Eq. 60, we can observe that for all $x\in\mathcal{X}_{<},a=\pi_{e}(x)$, the set $\mathcal{S}_{a}(x)$ is non-empty by definition of $\mathcal{X}_{<}$, and moreover that $m_{e}\geq\min(\mathbf{M}_{a}(x))$. As a result, we have it that $f_{Y}^{L}(x,a,m_{e},\epsilon_{Y})=F^{-1}_{Y\mid x,a,h_{\text{prev}}(m_{e})}(\epsilon_{Y})\sim P(Y\mid x,a,h_{\text{prev}}(m_{e}))$, and thus this term can be re-written as
    $\mathbf{1}\{X\in\mathcal{X}_{<},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\tilde{\mathcal{S}}_{<}(X)]$ (65)
    since for any $D\in\tilde{\mathcal{S}}_{<}(X)$, we have it that $\pi_{D}(X)=\pi_{e}(X)$ and $f_{M}(\pi_{D})=h_{\text{prev}}(m_{e})$ by definition of $h_{\text{prev}}(m_{e})$ and $\tilde{\mathcal{S}}_{<}$.
  • For Eq. 61, we can observe that for all $x\in\mathcal{X}_{\not\leq},a=\pi_{e}(x)$, the set $\mathcal{S}_{a}(x)$ is either empty, or non-empty where $f_{M}(\pi_{e})<\min(\mathbf{M}_{a}(x))$, by definition of $\mathcal{X}_{\not\leq}$. In either case, we have it that $f_{Y}^{L}=Y_{\text{min}}$, and so this term is equal to
    $\mathbf{1}\{X\in\mathcal{X}_{\not\leq},\pi_{e}(X)\neq a_{0}\}Y_{\text{min}}$ (66)
  • For Eq. 62, we can observe that for all $x\in\mathcal{X}_{*},a=\pi_{e}(x)=a_{0}$, the set $\mathcal{S}_{a}(x)$ is non-empty by definition of $\mathcal{X}_{*}$, and moreover that $f_{Y}^{L}$ is invariant to the choice of $m$, and so (by a similar argument to Eq. 59, this term is equal to $\mathbf{1}\{X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e}(X),M=f_{M}(\pi_{e})]$, which is equal to
    $\mathbf{1}\{X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{*}(X)]$ (67)
    from Eq. 23 of Section B.1
  • For Eq. 63, we can observe that for all $x\not\in\mathcal{X}_{*}$, the set $\mathcal{S}_{a}(x)$ is empty by definition of $\mathcal{X}_{*}$, and so $f_{Y}^{L}=Y_{\text{min}}$, and so this term is equal to
    $\mathbf{1}\{X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}Y_{\text{min}}$ (68)
Collecting terms Eqs. 64, 65, 66, 67 and 68 gives us that
@@
$\quad=\mathbb{E}\big{[}\mathbf{1}\{X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{=}(X)]$
$\qquad+\mathbf{1}\{X\in\mathcal{X}_{<},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\tilde{S}_{<}(X)]$
$\qquad+\mathbf{1}\{X\in\mathcal{X}_{\not\leq},\pi_{e}(X)\neq a_{0}\}Y_{\text{min}}$
$\qquad+\mathbf{1}\{X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{*}(X)]$
$\qquad+\mathbf{1}\{X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}Y_{\text{min}}\big{]}$
which is equivalent to $L(\pi_{e})$ using $\tilde{\mathcal{S}}_{<}$, and which completes the proof for the lower bound. For the upper bound, the argument is similar, and roughly symmetric, but uses the partition given by Eq. 36.
@@ Conclusion
Because it is possible to construct structural causal models that are consistent with our assumptions and that have counterfactual policy values that are exactly $L(\pi_{e})$ and $U(\pi_{e})$, the bounds in Assumption 3.3 cannot be improved without further assumptions. ∎
@@
Lemma B.1.
Let $\mathcal{S}^{\prime}(X)$ be a function that maps from $\mathcal{X}$ to any subset (including the empty set) of $\mathcal{D}$, and let $\mathcal{X}^{\prime}$ be a subset of $\mathcal{X}$. If $P(D\in\mathcal{S}^{\prime}(x)))>0,\forall x\in\mathcal{X}^{\prime}$, then
$\mathbb{E}\left[Y\frac{\mathbf{1}\{D\in\mathcal{S}^{\prime}(X)\}}{P(D\in\mathcal{S}^{\prime}(X))}\mathbf{1}\{X\in\mathcal{X}^{\prime}\}\right]=\mathbb{E}[\mathbb{E}[Y\mid D\in\mathcal{S}^{\prime}(X),X]\mathbf{1}\{X\in\mathcal{X}^{\prime}\}]$ (69)
@@
$\mathbb{E}\left[Y\frac{\mathbf{1}\{D\in\mathcal{S}^{\prime}(X)\}}{P(D\in\mathcal{S}^{\prime}(X)}\mathbf{1}\{X\in\mathcal{X}^{\prime}\}\right]$ $=\mathbb{E}\left[\mathbb{E}[Y\mathbf{1}\{D\in\mathcal{S}^{\prime}(X)\}\mid X]\frac{\mathbf{1}\{X\in\mathcal{X}^{\prime}\}}{P(D\in\mathcal{S}^{\prime}(X))}\right]$ (70)
$=\mathbb{E}\left[\mathbb{E}[Y\mid D\in\mathcal{S}^{\prime}(X),X]P(D\in\mathcal{S}^{\prime}(X)\mid X)\frac{\mathbf{1}\{X\in\mathcal{X}^{\prime}\}}{P(D\in\mathcal{S}^{\prime}(X))}\right]$ (71)
$=\mathbb{E}[\mathbb{E}[Y\mid D\in\mathcal{S}^{\prime}(X),X]\mathbf{1}\{X\in\mathcal{X}^{\prime}\}]$ (72)
where the first equality is well-defined on both sides by the assumption that for any $X\in\mathcal{X}^{\prime}$, $P(D\in\mathcal{S}^{\prime}(X))>0$. For the second-to-last line, note that this follows from the basic fact that $A,B,C$
@@
and the last line follows from the fact that $D\perp\!\!\!\perp X$ under Assumption 2.1, so that $P(D\in\mathcal{S}^{\prime}(X)\mid X)=P(D\in\mathcal{S}^{\prime}(X))$. ∎
@@
Note that Lemma B.1 applies to all of the pairs (e.g., $\mathcal{S}_{=}(X)$ and $X\in\mathcal{X}_{=}\cap\pi_{e}(X)\neq a_{0})$) used in Assumption 3.3. Thus, we can directly write the following through linearity of expectations and three applications of Lemma B.1.
$L(\pi_{e})$ $=\mathbb{E}[\mathbf{1}\{x\in\mathcal{X}_{=}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{=}(x)]$
$\quad+\mathbf{1}\{x\in\mathcal{X}_{<}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{<}(x)]$
$\quad+\mathbf{1}\{x\in\mathcal{X}_{\not\leq}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{min}$
$\quad+\mathbf{1}\{x\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{*}(x)]$
$\quad+\mathbf{1}\{x\not\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}Y_{min}]$
$=\mathbb{E}[Y\frac{\mathbf{1}\{D\in\mathcal{S}_{=}(x)\}}{P(D\in\mathcal{S}_{=}(x))}\mathbf{1}\{X\in\mathcal{X}_{=}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}$
$\quad+Y\frac{\mathbf{1}\{D\in\mathcal{S}_{<}(x)\}}{P(D\in S_{<}(x))}\mathbf{1}\{X\in\mathcal{X}_{<}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}$
$\quad+Y_{min}\mathbf{1}\{X\in\mathcal{X}_{\not\leq}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}$
$\quad+Y\frac{\mathbf{1}\{D\in S_{*}(x)\}}{P(D\in S_{*}(x))}\mathbf{1}\{X\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}$
$\quad+Y_{min}\mathbf{1}\{X\not\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}]$
@@
Because the sets $(X\in\mathcal{X}_{=}\cap\pi_{e}\neq a_{0})$, $(X\in\mathcal{X}_{<}\cap\pi_{e}\neq a_{0})$, $(X\in\mathcal{X}_{\not\leq}\cap\pi_{e}\neq a_{0})$, $(X\in\mathcal{X}_{*}\cap\pi_{e}=a_{0})$, and $(X\not\in\mathcal{X}_{*}\cap\pi_{e}=a_{0})$ are disjoint, only one product of the indicator functions inside the expectation above ever evaluates to $1$. Thus, we can equivalently express the expression inside the expectation above as the piecewise function
$\psi_{L}(X,D,Y)$

### v2

@@ Appendix D Proofs
    $f_{Y}^{L}(x,a,m,\epsilon_{Y})$ =$=\begin{cases}Y_{\text{min}},&\ \text{if }{\bf\Pi}^{a}(x)=\varnothing,\\ Y_{\text{min}},&\ \text{if }{\bf\Pi}^{a}(x)\neq\varnothing,m<\min(\mathbf{M}_{a}(x))\\ F_{Y\mid x,a,h_{\text{prev}}(m)}^{-1}(\epsilon_{Y}),&\ \text{if }{\bf\Pi}^{a}(x)\neq\varnothing,m>\min(\mathbf{M}_{a}(x)),m\not\in\mathbf{M}_{a}(x),\\ F_{Y\mid x,a,m}^{-1}(\epsilon_{Y}),&\ \text{if }{\bf\Pi}^{a}(x)\neq\varnothing,m\in\mathbf{M}_{a}(x),\end{cases}$ (56)
    $f_{Y}^{U}(x,a,m,\epsilon_{Y})$ =$=\begin{cases}Y_{\text{max}},&\ \text{if }{\bf\Pi}^{a}(x)=\varnothing,\\ Y_{\text{max}},&\ \text{if }{\bf\Pi}^{a}(x)\neq\varnothing,m>\max(\mathbf{M}_{a}(x))\\ F_{Y\mid x,a,h_{\text{next}}(m)}^{-1}(\epsilon_{Y}),&\ \text{if }{\bf\Pi}^{a}(x)\neq\varnothing,m<\max(\mathbf{M}_{a}(x)),m\not\in\mathbf{M}_{a}(x),\\ F_{Y\mid x,a,m}^{-1}(\epsilon_{Y}),&\ \text{if }{\bf\Pi}^{a}(x)\neq\varnothing,m\in\mathbf{M}_{a}(x),\end{cases}$ (57)
    where $F^{-1}_{Y\mid x,a,m}$ is the inverse conditional CDF of $Y$ given $X,A,M$, derived from $P$, and where ${\bf\Pi}^{a}(x)\coloneqq\{\pi\in\Pi\mid\pi(x)=a\}$ and $\mathbf{M}_{a}(x)\coloneqq\{f_{M}(\pi):\pi\in{\bf\Pi}^{a}(x)\}$, as defined previously above.
@@
We have now defined the SCMs $\mathcal{M}_{L},\mathcal{M}_{U}$, and shown that these SCMs are both consistent with our assumptions. We will now briefly verify that both SCMs give rise to the same observed distribution $P(X,Y,A,M,\Pi,D)$, and then demonstrate that these SCMs achieve the upper and lower bounds that are given in Remark 3.1.
First, we have it by construction that both SCMs yield the observed distribution $P(X,Y,A,M,\Pi,D)$, so it remains to demonstrate that they agree with the observed distribution $P(Y\mid X,A,M,D,\Pi)$, which we can write equivalently as $P(Y\mid X,A,M)$, since $D,\Pi\perp\!\!\!\perp Y\mid X,A,M$ under our assumed data-generating process. Note that $P(Y\mid X,A,M,\Pi,D)$ is only well-defined for $x,a,m$ with positive density (if $X$ is continuous) or probability mass (if $X$ is discrete). Assuming that $P(x)>0$ for all $x\in\mathcal{X}$, we have constructed $f_{Y}^{L},f_{Y}^{U}$ to agree with $P(Y\mid x,a,m)$ for all $a,m$ where there exists a trialed policy $\pi$ that outputs $a=\pi(x)$ with performance $m=f_{M}(\pi)$. We note that for any other value of $a^{\prime},m^{\prime}$, we have it that $P(a^{\prime},m^{\prime}\mid x)=0$, and hence the entire set $(x,a^{\prime},m^{\prime})$ has zero density, and it is precisely on these never-observed subsets of inputs where $\mathcal{M}_{L},\mathcal{M}_{U}$ disagree.
@@
$\mathbb{E}_{\mathcal{M}_{L}}[Y(A=\pi_{e},M=f_{M}(\pi_{e}))]$ $=\mathbb{E}\big{[}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\}f_{Y}^{L}(X,\pi_{e}(X),f_{M}(\pi_{e}),\epsilon_{Y})$ (58)
$\qquad+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,\pi_{e}(X)\neq a_{0}\}f_{Y}^{L}(X,\pi_{e}(X),f_{M}(\pi_{e}),\epsilon_{Y})$ (59)
$\qquad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\}f_{Y}^{L}(X,\pi_{e}(X),f_{M}(\pi_{e}),\epsilon_{Y})$ (60)
$\qquad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(X)=a_{0}\}f_{Y}^{L}(X,\pi_{e}(X),f_{M}(\pi_{e}),\epsilon_{Y})]$ (61)
@@
  • For Eq. 58, we can observe that for all $x\in\mathcal{X}$ satisfying $\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,a=\pi_{e}(x)$, the set ${\bf\Pi}^{a}(x)$ is non-empty by definition of $\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)$, and moreover that $m_{e}\geq\min(\mathbf{M}_{a}(x))$. As a result, we have it that $f_{Y}^{L}(x,a,m_{e},\epsilon_{Y})=F^{-1}_{Y\mid x,a,h_{\text{prev}}(m_{e})}(\epsilon_{Y})\sim P(Y\mid x,a,h_{\text{prev}}(m_{e}))$, and thus this term can be re-written as
    $\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)]$ (62)
    since for any $\pi_{i}\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)$, we have it that $\pi_{i}(X)=\pi_{e}(X)$ and $f_{M}(\pi_{i})=h_{\text{prev}}(m_{e})$ by definition of $h_{\text{prev}}(m_{e})$ and $\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)$.
  • For Eq. 59, we can observe that for all $x\in\mathcal{X}$ satisfying $\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,a=\pi_{e}(x)$, the set ${\bf\Pi}^{a}(x)$ is either empty, or non-empty where $f_{M}(\pi_{e})<\min(\mathbf{M}_{a}(x))$, by definition of $\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)$. In either case, we have it that $f_{Y}^{L}=Y_{\text{min}}$, and so this term is equal to
    $\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,\pi_{e}(X)\neq a_{0}\}Y_{\text{min}}$ (63)
  • For Eq. 60, we can observe that for all $x\in\mathcal{X}$ satisfying $\mathbf{\Pi}^{e}(X)\neq\varnothing,a=\pi_{e}(x)=a_{0}$, the set ${\bf\Pi}^{a}(x)$ is non-empty by definition of $\mathbf{\Pi}^{e}(X)$, and moreover that $f_{Y}^{L}$ is invariant to the choice of $m$, and so this term is equal to $\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e}(X),M=f_{M}(\pi_{e})]$, which is equal to
    $\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,\Pi\in\mathbf{\Pi}^{e}(X)]$ (64)
    from Eq. 24 of Section D.1
  • For Eq. 61, we can observe that for all $x\in\mathcal{X}$ satisfying $\mathbf{\Pi}^{e}(X)=\varnothing$, the set ${\bf\Pi}^{a}(x)$ is empty by definition of $\mathbf{\Pi}^{e}(X)$, and so $f_{Y}^{L}=Y_{\text{min}}$. Thus, this term is equal to
    $\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(X)=a_{0}\}Y_{\text{min}}$ (65)
Collecting terms Eqs. 62, 63, 64 and 65 gives us that
@@
$\quad=\mathbb{E}\big{[}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)]$
$\qquad+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,\pi_{e}(X)\neq a_{0}\}Y_{\text{min}}$
$\qquad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,\Pi\in\mathbf{\Pi}^{e}(X)]$
$\qquad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(X)=a_{0}\}Y_{\text{min}}\big{]}$
which is equivalent to $L(\pi_{e})$ and completes the proof for the lower bound. For the upper bound, the argument is similar, and roughly symmetric, but uses the partition given by Eq. 37.
@@ Conclusion
Because it is possible to construct structural causal models that are consistent with our assumptions and that have counterfactual policy values that are exactly $L(\pi_{e})$ and $U(\pi_{e})$, the bounds in Remark 3.1 cannot be improved without further assumptions. ∎
@@
First, note that the conditional expectation of $Y$ is always finite due to Assumption 3.3.
Lemma D.1.
Let ${\bf\Pi}^{\prime}(X)$ be a function that maps from $\mathcal{X}$ to any subset (including the empty set) of $\Pi$, and let $\mathcal{X}^{\prime}$ be a subset of $\mathcal{X}$. If $P(\Pi\in{\bf\Pi}^{\prime}(x))>0,\forall x\in\mathcal{X}^{\prime}$, then
$\mathbb{E}\left[Y\frac{\mathbf{1}\{\Pi\in{\bf\Pi}^{\prime}(X)\}}{P(\Pi\in{\bf\Pi}^{\prime}(X))}\mathbf{1}\{X\in\mathcal{X}^{\prime}\}\right]=\mathbb{E}[\mathbb{E}[Y\mid\Pi\in{\bf\Pi}^{\prime}(X),X]\mathbf{1}\{X\in\mathcal{X}^{\prime}\}]$ (66)
@@
$\mathbb{E}\left[Y\frac{\mathbf{1}\{\Pi\in{\bf\Pi}^{\prime}(X)\}}{P(\Pi\in{\bf\Pi}^{\prime}(X))}\mathbf{1}\{X\in\mathcal{X}^{\prime}\}\right]$ $=\mathbb{E}\left[\mathbb{E}[Y\mathbf{1}\{\Pi\in{\bf\Pi}^{\prime}(X)\}\mid X]\frac{\mathbf{1}\{X\in\mathcal{X}^{\prime}\}}{P(\Pi\in{\bf\Pi}^{\prime}(X))}\right]$ (67)
$=\mathbb{E}\left[\mathbb{E}[Y\mid\Pi\in{\bf\Pi}^{\prime}(X),X]P(\Pi\in{\bf\Pi}^{\prime}(X)\mid X)\frac{\mathbf{1}\{X\in\mathcal{X}^{\prime}\}}{P(\Pi\in{\bf\Pi}^{\prime}(X))}\right]$ (68)
$=\mathbb{E}[\mathbb{E}[Y\mid\Pi\in{\bf\Pi}^{\prime}(X),X]\mathbf{1}\{X\in\mathcal{X}^{\prime}\}]$ (69)
where the first equality is well-defined on both sides by the assumption that for any $X\in\mathcal{X}^{\prime}$, $P(\Pi\in{\bf\Pi}^{\prime}(X))>0$. For the second-to-last line, note that this follows from the basic fact that $A,B,C$
@@
and the last line follows from the fact that $\Pi\perp\!\!\!\perp X$ under Assumption 2.1, so that $P(\Pi\in{\bf\Pi}^{\prime}(X)\mid X)=P(\Pi\in{\bf\Pi}^{\prime}(X))$. ∎
@@
Note that Lemma D.1 applies to all of the pairs (e.g., $\mathbf{\Pi}^{e}(X)\neq\varnothing$ and $\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\}$) used in Remark 3.1. Thus, we can directly write the following through linearity of expectations and two applications of Lemma D.1.
$L(\pi_{e})$ $=\mathbb{E}[\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)]$
$\quad+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{min}$
$\quad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X,\Pi\in\mathbf{\Pi}^{e}(X)]$
$\quad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\}Y_{min}]$
$=\mathbb{E}[Y\frac{\mathbf{1}\{\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\}}{P(\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X))}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}$
$\quad+Y_{min}\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}$
$\quad+\frac{\mathbf{1}\{\Pi\in\mathbf{\Pi}^{e}(X)\}}{P(\Pi\in\mathbf{\Pi}^{e}(X))}\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\}$
$\quad+Y_{min}\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\}]$
We apply Lemma D.1 in two instances: one where the indicator function $\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}$ is turned on and one where the indicator function $\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing\}$ is turned on. In the case where $\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}$ is true, $P(\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(X))>0$ is also true because there is at least one trial policy in the set $\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)$. Similarly, when $\{\mathbf{\Pi}^{e}(X)\neq\varnothing\}$ is satisfied, $P(\Pi\in\mathbf{\Pi}^{e}(X))>0$ will also be satisfied. Thus, when applying Lemma D.1, the assumption required in the lemma that $P(\Pi\in{\bf\Pi}^{\prime}(x))>0$ is satisfied by the decomposition of the lower bound, and we do not require any additional assumptions regarding the probability of deploying a particular set of trial models.
@@
Because the sets $(\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\pi_{e}\neq a_{0})$, $(\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,\pi_{e}\neq a_{0})$, $(\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}=a_{0})$, and $(\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}=a_{0})$ are disjoint, only one product of the indicator functions inside the expectation above ever evaluates to $1$. Thus, we can equivalently express the expression inside the expectation above as the piecewise function
$\psi_{L}(Y,X,\Pi)$